OpenStudy (e.cociuba):

True or False. If F(x) is an antiderivative of f(x) and c is any constant, then F(x) + c is also an antiderivative of f(x).

OpenStudy (e.cociuba):

It's true?

OpenStudy (anonymous):

True:D

OpenStudy (anonymous):

ya its true

OpenStudy (e.cociuba):

Can u explain y?

OpenStudy (anonymous):

yupp ur right

OpenStudy (anonymous):

true is the answer

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Differentiate F(x) + c

OpenStudy (e.cociuba):

i new it thanks! :))

OpenStudy (anonymous):

oh no here comes JIM ... lol

jimthompson5910 (jim_thompson5910):

F(x) is the antiderivative of f(x) so by definition, d/dx[ F(x) ] = f(x)

OpenStudy (e.cociuba):

kk thanks guys! :))

jimthompson5910 (jim_thompson5910):

so d/dx[ F(x) + C ] = d/dx[ F(x) ] + d/dx[ C ] d/dx[ F(x) + C ] = d/dx[ F(x) ] + 0 ... derivative of a constant is 0 d/dx[ F(x) + C ] = f(x)

OpenStudy (anonymous):

oh no good luck reading tht lol

jimthompson5910 (jim_thompson5910):

lol it's not too bad actually

OpenStudy (e.mccormick):

The infinite number of antiderivatives all only vary by a constant. Take a curve an curve, and shift the y axis up or down. Does that change the derivative at all? No. Thus, the antiderivative +c is related to this.

OpenStudy (anonymous):

yeah well maybe not for u...

OpenStudy (e.cociuba):

No its fine, thanks guys lol :)

jimthompson5910 (jim_thompson5910):

but the whole thing to take away from this is that integral of f(x) = F(x) + C which is what a lot of people (including me) forget...that +C

OpenStudy (anonymous):

Ur Welcome:) @e.cociuba

OpenStudy (e.cociuba):

Yea I tend to forget and it confuses me sometimes lol

OpenStudy (e.cociuba):

kk thanks @jim_thompson5910 @LizzyLove<3 and @e.mccormick :))

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Peace Yall ;P